# Benchmarking Language Model Creativity: A Case Study on Code Generation

### v1

@@ 4 State-Aware and Human-Grounded Evaluation of Machine Creativity
Given NeoCoder dataset, we introduce our metric of creativity NeoGauge for a given model $\mathbf{G}_{\text{LM}}$. Our NeoCoder at state $t$ ($t\leq T$) is: $\mathcal{D}_{t}=\{(x^{i},\mathcal{C}_{t}^{i}=\{\tau^{i}_{1},\tau^{i}_{2},\cdots,\tau^{i}_{t}\})\mid i=1,2,\cdots,n\}$, where $i$ is the problem index. To evaluate the creativity of the testing model $\mathbf{G}_{\text{LM}}$ [^5] We use distinct notations for the testing language model, $\mathbf{G}_{\text{LM}}$, and the augmentation language model, $\mathbf{P}_{\text{LM}}$, to highlight their different roles. at state $t$, we feed $\mathcal{D}_{t}$ to $\mathbf{G}_{\text{LM}}$ to obtain its predictions:
$\mathcal{Y}_{t}=\{y^{i}_{t}\sim\mathbf{G}_{\text{LM}}(x^{i}\oplus\mathcal{C}^{i}_{t})\mid|\mathcal{C}_{t}^{i}|=t,\forall(x^{i},\mathcal{C}_{t}^{i})\in\mathcal{D}_{t}\},$ (1)
where $|\mathcal{C}_{t}^{i}|$ denotes the cardinality of constraints set. It ensures that at any given state $t$, the questions we evaluated always have $t$ distinct constraints. Below, we present how we compute convergent and divergent creativity and introduce NeoGauge metric that unifies them.
Figure 3: Example of NeoGauge computation. The question comes from our NeoCoder dataset with ID 1829B and testing model $\mathbf{G}_{\text{LM}}$ here is GPT-4. For each state, we compute NeoGauge (Eq.4) as the probability of LM generating correct solutions that meet the given constraints (convergent creativity defined in Eq.2) and also exhibit $\mathcal{H}$-creativity (divergent creativity defined in Eq.3). However, none of the above three solutions are considered to be “creative” since convergent solutions may lack divergent creativity (e.g., state $t=0$). Alternatively, LMs’ hallucinated responses resulting in high $\mathcal{H}$-creativity, but often lack correctness and constraint following (e.g., state $t=1$). Therefore, truly creative works should not only be innovative but also appropriately solve a problem.
@@
To evaluate $\mathbf{G}_{\text{LM}}$’s convergent thinking ability, we examine two characteristics of generated solutions: whether they are correct and follow the given constraints. Therefore, given $\mathcal{Y}_{t}$ from Eq.1, we define its convergent creativity as follows:
$\textbf{convergent}(\mathbf{G}_{\text{LM}},t)$ $=\frac{1}{|\mathcal{Y}_{t}|}\sum_{y^{i}_{t}\in\mathcal{Y}_{t}}\mathbbm{1}^{\mathcal{T}^{i}_{t}\cap\mathcal{C}^{i}_{t}=\varnothing}\times\mathbbm{1}^{\text{Correct}(y_{t}^{i})},\;\text{where}\;\mathcal{T}^{i}_{t}\sim\mathbf{P}_{\text{LM}}(y^{i}_{t}).$ (2)
$\mathbbm{1}^{\text{Correct}(y^{i}_{t})}$ is a measure of program correctness, set to 1 if the generated solution passes all test examples, otherwise it is 0. We use the augmentation model $\mathbf{P}_{\text{LM}}$ to detect all atomic techniques $\mathcal{T}^{i}_{t}$ used in solution $y^{i}_{t}$, and compare them with the given constraint list $\mathcal{C}^{i}_{t}$ to check if the solution follows the given constraints. In Fig.3 examples, only the solution generated at $t=0$ (which does not involve any constraint) exhibits convergent creativity.
@@
As discussed earlier in §2, the primary focus of our evaluation is on $\mathcal{H}$-creativity, which requires a juxtaposition of model solutions with historical human solutions. Let’s consider a finite set of correct human written solutions with size $m$, denoted as $\mathcal{H}^{i}$, for problem $x^{i}$. Rather than directly comparing solutions using certain sentence-level similarity scores, as done by few prior works such as DeLorenzo et al. [16], we break down the comparison to the atomic technique level, which is more interpretable and generalizable across varying solutions. Our divergent creativity score is defined as:
$\textbf{divergent}(\mathbf{G}_{\text{LM}},t)=\frac{1}{|\mathcal{Y}_{t}|}\sum_{y^{i}_{t}\in\mathcal{Y}_{t}}\frac{|\mathcal{T}^{i}_{t}\setminus\widehat{\mathcal{T}}^{i}|}{|\mathcal{T}^{i}_{t}|},$
$\text{where}\;\mathcal{T}^{i}_{t}\sim\mathbf{P}_{\text{LM}}(y^{i}_{t}),\;\widehat{\mathcal{T}}^{i}=\bigcup_{j=1}^{m}\widehat{\mathcal{T}}^{i}_{j}\sim\mathbf{P}_{\text{LM}}(\hat{y}^{i}_{j}),\;\hat{y}^{i}_{j}\in\mathcal{H}^{i}.$ (3)
Here we first find all atomic techniques $\widehat{\mathcal{T}}^{i}$ used by $m$ human solutions and atomic techniques used in the model solution $\mathcal{T}^{i}_{t}$ at state $t$. We then compute the $\mathcal{H}$-creativity as the ratio of techniques used by $\mathbf{G}_{\text{LM}}$ that have never been used in the human solution set. For example, as shown in Fig.3 at state $t=1$, among the three techniques identified within the generated solution, only the recursion has never been used by humans, thereby resulting in a ratio of $\frac{1}{3}$. Finally, we average ratios across different problems to obtain the final $\mathcal{H}$-creativity at state $t$.
@@
Given the above definitions, NeoGauge of $\mathbf{G}_{\text{LM}}$ at state t can be formalized as:
@@
$\text{where}\;\mathcal{Y}_{t}=\{y^{i}_{t}\sim\mathbf{G}_{\text{LM}}(x^{i}\oplus\mathcal{C}^{i}_{t})\mid|\mathcal{C}_{t}^{i}|=t,\forall(x^{i},\mathcal{C}_{t}^{i})\in\mathcal{D}_{t}\},$
$\mathcal{T}^{i}_{t}\sim\mathbf{P}_{\text{LM}}(y^{i}_{t}),\;\widehat{\mathcal{T}}^{i}=\bigcup_{j=1}^{m}\widehat{\mathcal{T}}^{i}_{j}\sim\mathbf{P}_{\text{LM}}(\hat{y}^{i}_{j}),\;\hat{y}^{i}_{j}\in\mathcal{H}^{i}.$ (4)
@@
$\textbf{convergent}(\mathbf{G}_{\text{LM}},t)$ Convergent creativity of $\mathbf{G}_{\text{LM}}$ at state $t$ Eq.2 Tbl.4, Fig.6, 8
$\textbf{divergent}(\mathbf{G}_{\text{LM}},t)$ Divergent creativity of $\mathbf{G}_{\text{LM}}$ at state $t$ Eq.3 Tbl.4, Fig.6, 8
NeoGauge@t Creativity evaluation of $\textbf{G}_{\text{LM}}$ at state $t$ Eq.4 Tbl.4, Fig.4
@@
convergent(human, t) convergent creativity of human at state $t$ Eq.5 Fig.6
divergent(human) divergent creativity of human at state 00 Eq.6 Fig.6
@@ 5 Experiments and Results
We benchmark language model (§5.2) and evaluate reasoning strategies (§5.3) for creativity.
@@ 5.1 Experimental Setup
We use GPT-4 as the augmentation model $\mathbf{P}_{\text{LM}}$. We benchmark the creativity performance of the following testing models $\mathbf{G}_{\text{LM}}$: GPT-4 [37], GPT-3.5 [38], Claude 3 Sonnet (Claude-3) [4], Llama3-70B [1], Llama2-70B [54], CodeLlama-34B-Python (CodeLlama-34B) [43], CodeGemma-7B [21], and Mistral-7B [28]. We access all non-proprietary models through Huggingface Transformers [55].
@@
Beyond the three proposed metrics for evaluating convergent, divergent and overall creativity, we also compute pass@1 [12] and constraint following ratio for further comparison in Table 4. NeoGauge@T actually is a joint probability of $\mathbf{G}_{\text{LM}}$ being both convergent and divergent creative at state $t$. Therefore, we also report the cumulative NeoGauge across states in Fig.4, which indicates the model’s maximum creativity performance boundary. Additionally, we compute human convergent and divergent creativity in Fig.6 to compare LLM with human creativity performance (details in Appendix A). We summarize all used metrics in Table 3.
@@ 5.2 Benchmarking Language Model Creativity
A number of psychological investigators have studied the link between creativity and intelligence [25], agreeing on two key points: (1) creative individuals tend to have higher intelligence [41], and (2) people with extremely high intelligence not necessarily to be extremely creative [17]. We re-examine the two findings on LLMs and answer: Are larger LLMs more creative? Do extremely large models of equal size exhibit comparable creativity? Our investigation is based on the widely accepted hypothesis that language model size correlates positively with intelligence [30, 32, 62].
@@
GPT-4 is the most creative LLM thus far.
We visualize NeoGauge and cumulative NeoGauge in Fig.4. GPT-4 consistently has the highest NeoGauge almost at every state $t$. While others (e.g., Claude-3 and Llama3-70B) have a close NeoGauge@0 score to GPT-4, their NeoGauge quickly decreases to 0 within the next 2 states. According to cumulative NeoGauge, GPT-4 also has the highest creativity performance boundary, followed by Claude-3 and Llama3-70B, greatly outperforming smaller models such as GPT-3.5 and Llama2-70B. These observations could potentially answer the above two questions: larger LLMs are generally more creaitive, but extremely large LLM is not necessarily exhibiting extremely creative performance. In Fig.5, we provide example outputs from each model to show their different creativity abilities.
Figure 6: A comparison of LLM and human creativity. //// denotes the performance difference of convergent creativity, and \\\\ denotes the difference of divergent creativity. Current LLMs still hardly demonstrate human-like creativity.
LLM creativity is still far behind humans.
Fig.6 displays the comparison between LLM and human creativity evaluation. LLM shows slightly better performance in divergent creativity compared to humans. However, human divergent creativity, as computed from Eq.6, is a constant and only reflects the initial divergent creativity at state $t=0$. The performance gap between them in higher states remains undetermined. Additionally, Both human and LLM convergent creativity declines drastically over the increase in state $t$, which follows our expectation that there is a trade-off between solution quality and novelty. When stress-testing humans or LLMs to look for more creative solutions, they are very likely to make mistakes and may copy previous solutions during the process. However, LLM convergent creative performance is much worse than humans. Future works should focus on how to narrow the gap (//// area) between human and LLM convergent creativity.
@@ 5.3 Evaluating Reasoning Strategies for Creativity
We evaluate four reasoning strategies on our NeoCoder dataset to further study the correlation between augmented machine intelligence and creativity: Whether such intelligence-enhancing techniques also improve creative thinking? We implement the following four works that are specifically designed for programming tasks with different reasoning strategies:
  • MCTS: Zhang et al. [59] propose a novel decoding method that uses Monte-Carlo tree search (MCTS) to generate better programs using the pass rate as reward.
  • Self-Correction: Shinn et al. [45] use verbal feedback from a reflection agent to reinforce the performance of an actor agent in code generation.
  • Planning: Jiang et al. [29] design a planning module to let LLM plan out concise solution steps from the intent, followed by an implementation module to generate code step by step.
  • Sampling: Chen et al. [12] generate $k$ samples and compute the probability that at least one of the $k$-generated code samples for a problem passes the unit tests. For creativity evaluation, we generate $k=5$ samples for each problem and report the NeoGauge from samples that have the highest convergent and divergent creativity, $\mathbbm{1}^{\mathcal{T}^{i}_{t}\cap\mathcal{C}^{i}_{t}=\varnothing}\times\mathbbm{1}^{\text{Correct}(y_{t}^{i})}\times\frac{|\mathcal{T}^{i}_{t}\setminus\widehat{\mathcal{T}}^{i}|}{|\mathcal{T}^{i}_{t}|}$ in Eq.4, among $k=5$ samples.
Note that these methods are originally applicable to different kinds of models. Considering the computation complexity and the cost, we re-evaluate MCTS on the open-source language model (CodeGemma-7B [21]) and re-evaluate others on the proprietary model (GPT-3.5 [38]).
@@
According to Fig.7, all reasoning strategies except sampling help to improve the model’s convergent creativity thinking ability on multiple states, as they are fundamentally designed to improve the accuracy. Conversely, only MCTS successfully enhances divergent creativity, due to it rolling out numerous paths during the expansion. Strategies like self-correction, planning, and sampling, which operate on a single trial or path, fail to explore divergent solutions.

### v2

@@ 4 State-Aware and Human-Grounded Evaluation of Machine Creativity
Augmentation model vs target model.
So far, we have used $\mathbf{P}_{\text{LM}}(.)$ to denote the augmentation model, the language model used for dataset construction and extracting atomic techniques. Here, we introduce $\mathbf{G}_{\text{LM}}(\cdot)$ to represent the target language model, whose creativity we evaluate using our dataset and the augmentation model $\mathbf{P}_{\text{LM}}(\cdot)$.
Setup.
Here we introduce our metric of creativity NeoGauge for a given model $\mathbf{G}_{\text{LM}}$ and given NeoCoder. Denote instances of NeoCoder at state $t$ ($t\leq T$) as:
$\mathcal{D}_{t}=\Big{\{}(x^{i},\mathcal{C}_{t}^{i}=\{\tau^{i}_{1},\tau^{i}_{2},\cdots,\tau^{i}_{t}\})\Big{\}}_{i=1}^{n},$
where $i$ is the problem index. To evaluate the creativity of the testing model $\mathbf{G}_{\text{LM}}$ at state $t$, we feed $\mathcal{D}_{t}$ to $\mathbf{G}_{\text{LM}}$ to obtain its predictions:
$\mathcal{Y}_{t}=\Big{\{}y^{i}_{t}\sim\mathbf{G}_{\text{LM}}(x^{i}\oplus\mathcal{C}^{i}_{t})$ $\Bigm{|}|\mathcal{C}_{t}^{i}|=t,$
$\forall(x^{i},\mathcal{C}_{t}^{i})\in\mathcal{D}_{t}\Big{\}}.$ (1)
Here $|\mathcal{C}_{t}^{i}|$ denotes the cardinality of the constraints set. The constraint $|\mathcal{C}_{t}^{i}|=t$ ensures that at a given state $t$, the questions we evaluated always have $t$ distinct constraints. Below, we present how we compute convergent and divergent creativity and introduce NeoGauge metric that unifies them.
Figure 3: Example of NeoGauge computation. The question comes from our NeoCoder dataset with ID 1829B and testing model $\mathbf{G}_{\text{LM}}$ here is GPT-4. For each state, we compute NeoGauge (Eq.4) as the probability of LM generating correct solutions that meet the given constraints (convergent creativity defined in Eq.2) and also exhibit $\mathcal{H}$-creativity (divergent creativity defined in Eq.3). However, none of the above three solutions are considered to be “creative” since convergent solutions may lack divergent creativity (e.g., state $t=0$). Alternatively, LLMs’ hallucinated responses resulting in high $\mathcal{H}$-creativity, but often lack correctness and constraint following (e.g., state $t=1$). Therefore, truly creative works should not only be innovative but also appropriately solve a problem.
@@
To evaluate $\mathbf{G}_{\text{LM}}$’s convergent thinking ability, we examine two characteristics of generated solutions: whether they are correct and whether they follow the given constraints. Therefore, given $\mathcal{Y}_{t}$ from Eq.1, we define its convergent creativity as follows:
$\textbf{convergent}(\mathbf{G}_{\text{LM}},t)=$
$\hskip 45.52458pt\frac{1}{|\mathcal{Y}_{t}|}\sum_{y^{i}_{t}\in\mathcal{Y}_{t}}\mathbbm{1}^{\mathcal{T}^{i}_{t}\cap\mathcal{C}^{i}_{t}=\varnothing}\times\mathbbm{1}^{\text{Correct}(y_{t}^{i})},$ (2)
where atomic techniques $\mathcal{T}^{i}_{t}\sim\mathbf{P}_{\text{LM}}(y^{i}_{t})$. $\mathbbm{1}^{\text{Correct}(y^{i}_{t})}$ is a measure of program correctness, set to 1 if the generated solution passes all the test examples. Otherwise it is 0. We use the augmentation model $\mathbf{P}_{\text{LM}}$ to detect all atomic techniques $\mathcal{T}^{i}_{t}$ used in solution $y^{i}_{t}$, and compare them with the given constraint list $\mathcal{C}^{i}_{t}$ to check if the solution follows the given constraints. In Figure 3 examples, only the solution generated at $t=0$ (which does not involve any constraint) exhibits convergent creativity.
@@
As discussed earlier in §2, a primary focus of our evaluation is on $\mathcal{H}$-creativity, which requires a juxtaposition of model solutions with historical human solutions. Let’s consider a finite set of correct human written solutions with size $m$, denoted as $\mathcal{H}^{i}$, for problem $x^{i}$. Rather than directly comparing solutions using certain sentence-level similarity scores, as done by a few prior works such as DeLorenzo et al. (2024), we break down the comparison to the atomic technique level, which is more interpretable and generalizable across varying solutions. Our divergent creativity score is defined as:
$\textbf{divergent}(\mathbf{G}_{\text{LM}},t)=\frac{1}{|\mathcal{Y}_{t}|}\sum_{y^{i}_{t}\in\mathcal{Y}_{t}}\frac{|\mathcal{T}^{i}_{t}\setminus\widehat{\mathcal{T}}^{i}|}{|\mathcal{T}^{i}_{t}|},$ (3)
where $\mathcal{T}^{i}_{t}\sim\mathbf{P}_{\text{LM}}(y^{i}_{t})$ are the atomic techniques used in the model solutions, and $\widehat{\mathcal{T}}^{i}$ indicate all the atomic techniques used by $m$ human solutions, defined as: $\widehat{\mathcal{T}}^{i}=\bigcup_{j=1}^{m}\big{\{}\widehat{\mathcal{T}}^{i}_{j}\sim\mathbf{P}_{\text{LM}}(\hat{y}^{i}_{j}),\;\hat{y}^{i}_{j}\in\mathcal{H}^{i}\big{\}}.$ We then compute the $\mathcal{H}$-creativity as the ratio of techniques used by $\mathbf{G}_{\text{LM}}$ that have never been used in the human solution set. For example, as shown in Figure 3 at state $t=1$, among the three techniques identified within the generated solution, only the recursion has never been used by humans, thereby resulting in a ratio of $\frac{1}{3}$. Finally, we average ratios across different problems to obtain the final $\mathcal{H}$-creativity at state $t$.
@@
Given the above definitions, NeoGauge of $\mathbf{G}_{\text{LM}}$ at state t can be formalized:
@@
$\frac{1}{|\mathcal{Y}_{t}|}\sum_{y^{i}_{t}\in\mathcal{Y}_{t}}\underbrace{\mathbbm{1}^{\mathcal{T}^{i}_{t}\cap\mathcal{C}^{i}_{t}=\varnothing}\mathbbm{1}^{\text{Correct}(y_{t}^{i})}}_{\text{Convergent Creativity}}\times\underbrace{\frac{|\mathcal{T}^{i}_{t}\setminus\widehat{\mathcal{T}}^{i}|}{|\mathcal{T}^{i}_{t}|}}_{\text{Divergent Creativity}},$ (4)
where $\mathcal{Y}_{t}=\{y^{i}_{t}\sim\mathbf{G}_{\text{LM}}(x^{i}\oplus\mathcal{C}^{i}_{t})\mid|\mathcal{C}_{t}^{i}|=t,\forall(x^{i},\mathcal{C}_{t}^{i})\in\mathcal{D}_{t}\}$ (defined in Eq.1), $\mathcal{T}^{i}_{t}\sim\mathbf{P}_{\text{LM}}(y^{i}_{t})$ (defined in Eq.2), $\widehat{\mathcal{T}}^{i}=\bigcup_{j=1}^{m}\big{\{}\widehat{\mathcal{T}}^{i}_{j}\sim\mathbf{P}_{\text{LM}}(\hat{y}^{i}_{j}),\;\hat{y}^{i}_{j}\in\mathcal{H}^{i}\}$ (defined in Eq.3).
@@
$\textbf{convergent}(\mathbf{G}_{\text{LM}},t)$ Convergent creativity of $\mathbf{G}_{\text{LM}}$ at state $t$ Eq.2 Table 3, Figure 5, 7
$\textbf{divergent}(\mathbf{G}_{\text{LM}},t)$ Divergent creativity of $\mathbf{G}_{\text{LM}}$ at state $t$ Eq.3 Table 3, Figure 5, 7
NeoGauge@t Creativity evaluation of $\textbf{G}_{\text{LM}}$ at state $t$ Eq.4 Table 3, Figure 4
@@
convergent(human, t) convergent creativity of human at state $t$ Eq.5 Figure 5
divergent(human) lowest divergent creativity of human at state 00 Eq.6 Figure 5
@@ 5 Experiments and Results
We report the creativity of current LLMs (§5.2) and evaluate different reasoning strategies (§5.3) for creativity.
@@ 5.1 Experimental Setup
We use GPT-4 as the augmentation model $\mathbf{P}_{\text{LM}}$. We benchmark the creativity performance of the following target models $\mathbf{G}_{\text{LM}}$: GPT-4 (OpenAI, 2024), GPT-3.5 (Ouyang et al., 2022), Claude 3 Sonnet (Claude-3) (Anthropic, 2024), Llama3-70B (AI@Meta, 2024), Llama2-70B (Touvron et al., 2023), CodeLlama-34B-Python (CodeLlama-34B) (Rozière et al., 2024), CodeGemma-7B (Google, 2024), and Mistral-7B (Jiang et al., 2023a). We access all non-proprietary models through Huggingface Transformers (Wolf et al., 2019). Following the parameter choice by Zhang et al. (2023), we apply a sampling temperature of 1 for code generation.
@@
Beyond the three proposed metrics for evaluating convergent, divergent and overall creativity, we also compute pass@1 (Chen et al., 2021) and constraint following ratio for further comparison in Table 3. NeoGauge@T actually is a joint probability of $\mathbf{G}_{\text{LM}}$ being both convergent and divergent creative at state $t$. Therefore, we also report the cumulative NeoGauge across states in Figure 4, which indicates the model’s maximum creativity performance boundary. Additionally, we compute human convergent and divergent creativity in Figure 5 to compare LLM with human creativity performance (details in Appendix B). We summarize all used metrics in Table 2.
@@ 5.2 Benchmarking Language Model Creativity
A number of psychological investigators have studied the link between creativity and intelligence (Holyoak and Morrison, 2005), agreeing on two key points: (1) creative individuals tend to have higher intelligence (Renzulli, 2005), and (2) people with extremely high intelligence not necessarily to be extremely creative (Faris et al., 1962). We re-examine the two findings on LLMs and answer: Are larger LLMs more creative? Do extremely large models of equal size exhibit comparable creativity? Our investigation is based on the widely accepted hypothesis that language model size correlates positively with intelligence (Kaplan et al., 2020; Liu et al., 2023; Zhao et al., 2023).
GPT-4 is the most creative LLM thus far.
We visualize NeoGauge and cumulative NeoGauge in Figure 4. GPT-4 consistently has the highest NeoGauge almost at every state $t$. While others (e.g., Claude-3 and Llama3-70B) have a close NeoGauge@0 score to GPT-4, their NeoGauge quickly decreases to 0 within the next two states. According to cumulative NeoGauge, GPT-4 also has the highest creativity performance boundary, followed by Claude-3 and Llama3-70B, greatly outperforming smaller models such as GPT-3.5 and Llama2-70B. These observations could potentially answer the above two questions: larger LLMs are generally more creaitive, but extremely large LLM is not necessarily exhibiting extremely creative performance. In Figure 9, we provide example outputs from each model to show their different creativity abilities.
@@
Figure 5: A comparison of LLM and human creativity. //// denotes the performance difference of convergent creativity, and \\\\ denotes the difference of divergent creativity. We observe that Current LLMs still hardly demonstrate human-like creativity.
Which is more creative: machine or human?
Figure 5 displays the creativity comparison between LLM and humans. LLM demonstrates minimally better performance in divergent creativity compared to humans at their lowest level (Eq.6). However, humans have significantly greater convergent creativity than LLMs in early states (prior to state 3). Thus, we reach a tentative conclusion that, in problem-solving settings, LLMs in Figure 5 barely exhibit human-like creativity. Future works could focus on measuring human divergent creativity across states to enable a fairer creativity comparison. Moreover, we observe that both human and LLM convergent creativity declines drastically over the increase in state $t$, which follows our expectation that there is a trade-off between solution quality and novelty. When stress-testing humans or LLMs to look for more creative solutions, they are very likely to make mistakes and may copy previous solutions during the process.
@@ 5.3 Evaluating Reasoning Strategies for Creativity
We evaluate four reasoning strategies on our NeoCoder dataset to further study the correlation between augmented machine intelligence and creativity: Whether such intelligence-enhancing techniques also improve creative thinking? We implement the following four works that are specifically designed for programming tasks:
  • MCTS: Zhang et al. (2023) propose a novel decoding method that uses Monte-Carlo Tree Search (MCTS) to generate better programs using the pass rate as reward.
  • Self-Correction: Shinn et al. (2023) use verbal feedback from a reflection agent to reinforce the performance of an agent in code generation.
  • Planning: Jiang et al. (2023b) design a planning module to let LLM plan out concise solution steps from the intent, followed by an implementation module to generate code step by step.
  • Sampling: Chen et al. (2021) generate $k$ samples and compute the probability that at least one of the $k$-generated code samples for a problem passes the unit tests. For creativity evaluation, we generate $k=5$ samples for each problem and report the NeoGauge from samples that have the highest convergent and divergent creativity, $\mathbbm{1}^{\mathcal{T}^{i}_{t}\cap\mathcal{C}^{i}_{t}=\varnothing}\times\mathbbm{1}^{\text{Correct}(y_{t}^{i})}\times\frac{|\mathcal{T}^{i}_{t}\setminus\widehat{\mathcal{T}}^{i}|}{|\mathcal{T}^{i}_{t}|}$ in Eq.4, among $k=5$ samples.
Note that these methods are originally applicable to different kinds of models. Considering the computation complexity and the cost, we re-evaluate MCTS on the open-source language model (CodeGemma-7B (Google, 2024)) and re-evaluate others on the proprietary model (GPT-3.5).
@@
According to Figure 6, all reasoning strategies except sampling help to improve the model’s convergent creativity thinking ability on multiple states, as they are fundamentally designed to improve the accuracy. Conversely, only MCTS successfully enhances divergent creativity, due to it rolling out numerous paths during the expansion. Strategies like self-correction, planning, and sampling, which operate on a single trial or path, fail to explore divergent solutions.